OpenStudy (anonymous):

Compound angle formula? how do i solve Tan 23pie/12

OpenStudy (anonymous):

So with all trigonometric functions, you want to think about how their value applies to the unit circle.

OpenStudy (anonymous):

For sin(x) and cos(x), \[0 \le y \le 1\]. Now, 'tangent' is just sin(x)/cos(x). So, in reality, your problem is: \[\sin(23\pi/12)/\cos(23\pi/12)\]

OpenStudy (anonymous):

Do you know how to apply the unit circle in trigonometry?

OpenStudy (anonymous):

In any case, usually you don't need to do sin (or cos) of (π/12) in any facet so you're probably okay leaving it like that.